# Learning Exposure Correction in Dynamic Scenes

### v2

@@ Abstract
Capturing videos with wrong exposure usually produces unsatisfactory visual effects. While image exposure correction is a popular topic, the video counterpart is less explored in the literature. Directly applying prior image-based methods to input videos often results in temporal incoherence with low visual quality. Existing research in this area is also limited by the lack of high-quality benchmark datasets. To address these issues, we construct the first real-world paired video dataset, including both underexposure and overexposure dynamic scenes. To achieve spatial alignment, we utilize two DSLR cameras and a beam splitter to simultaneously capture improper and normal exposure videos. In addition, we propose a Video Exposure Correction Network (VECNet) based on Retinex theory, which incorporates a two-stream illumination learning mechanism to enhance the overexposure and underexposure factors, respectively. The estimated multi-frame reflectance and dual-path illumination components are fused at both feature and image levels, leading to visually appealing results. Experimental results demonstrate that the proposed method outperforms existing image exposure correction and underexposed video enhancement methods. The code and dataset will be available soon.
@@ 1 Introduction
Figure 1: The upper part is the benchmark for our proposed video exposure correction. Taking underexposed video as an example, the lower part is the visual comparison with image-based method. LACT [2] takes single frames as input and results in temporal exposure incoherence with low visual quality. Our method utilizes temporal information to maintain the same exposure.
When taking photos or videos across various scenarios and lighting conditions, the results may not always be ideal due to low light, backlight, or intense illumination factors. Images and videos captured with underexposure or overexposure often yield unsatisfactory visual effects. When enhancing such mixed multiple exposure data, there are large discrepancies in the training exposure correction process of the network. This challenge is further compounded when dealing with videos captured in dynamic scenes, where cameras or objects undergo substantial movement. This paper focuses on correcting improperly exposed videos taken from dynamic scenes with suboptimal lighting conditions, which has not been well explored.
In recent years, the development of paired image datasets [3, 1] has spurred the creation of several methods [12, 11, 29, 23, 34] focused on learning image exposure correction through supervised learning. They learn a mapping relation from images taken in underexposure or overexposure conditions to those with normal exposures. However, these methods generally do not handle videos of dynamic scenes, as directly applying them in a frame-by-frame manner often results in temporal incoherence with low visual quality [19, 15], as shown in Fig. 1. A notable portion of some methods [31, 25, 33, 8, 4] is rooted in the Retinex theory [14] to achieve underexposed image or video enhancement. The objective involves modeling the relationship between the illumination and reflectance components through extensive data-driven learning. This cannot be directly applied to multiple exposure correction tasks with opposite optimization goals. To date, there is no current work exploring real-world video exposure correction datasets and methods. Therefore, we propose a benchmark for training and evaluating video multi-exposure correction.
@@
On the one hand, we aim to collect a high-quality video dataset with both underexposure and overexposure, and the main difficulties are as follows. First, collecting paired video datasets requires capturing two videos, one under abnormal exposure conditions and another under normal light of the same dynamic scene and camera motion. Second, it requires accurate alignment of each pair of corresponding frames in the two videos in both spatial and temporal dimensions. To overcome the above challenges, we design a two-camera system with a beam splitter to ensure no parallax between the two cameras with different aperture and ISO parameter settings. Two DSLR cameras simultaneously capture improper and normal exposure videos. Then we perform precise alignment on the captured pairs to make them aligned with each other. Furthermore, the well-exposed videos are manually rendered by professional photographers and then serve as the ground truth of the dataset. Finally, we construct a new dataset of 119 high-quality spatially-aligned videos named as DIME dataset, standing for “Dynamic scenes In Multiple Exposure”. It contains diverse real-world scenes, large camera and object motions, and accurate spatial alignment.
On the other hand, there are varying degrees of improper exposure between videos or between frames. We propose an end-to-end Video Exposure Correction Network (VECNet) to address both overexposure and underexposure problems in videos. Specifically, based on Retinex theory, we introduce the dual-stream illumination mechanism to construct illumination maps and enhanced images for the overexposure and underexposure branches. Then, to learn aligned reflectance components from consecutive frames, we introduce a novel multi-frame alignment strategy. Finally, a two-stage process is proposed to synthesize dual illumination and reflectance for restoration at feature and image levels. Our contributions are as follows:
@@
  • We construct the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multiple exposures, camera and object motions, and precise spatial alignment.
@@
  • We propose an effective exposure correction network based on Retinex theory to enhance overexposed and underexposed videos.
  • We conduct extensive experiments to demonstrate the superiority of our dataset and method.
@@ 2 Our DIME Dataset
Figure 2: We built the optical system to capture the over-/under- and normal exposure video pairs.
Figure 3: Statistics on our DIME dataset. (a) is the input-ground truth brightness mapping curve statistics. (b) Luminance distribution for overexposure videos. (c) Luminance distribution for underexposure videos. (d) Luminance distribution for ground truth videos.
To learn video exposure correction tasks by training our model, we need a large number of videos including realistic overexposure and underexposure errors and corresponding paired and properly exposed ground truth videos. As discussed in Sec. 2, such datasets are currently not publicly available to support video exposure correction research. Therefore, our first task is to create a novel dataset that captures various videos in dynamic and real-world scenes.
@@ 2.1 Hardware Design
Capturing multiple exposure image pairs for static scenes can be easily realized with a single exposure adjustable DSLR camera. However, For dynamic exposed video capturing, we need to utilize two cameras with different aperture and ISO parameter settings. To keep the cameras in sync, we use an infrared remote control to signal two camera-mounted receivers for simultaneous capture. To prevent parallax problems caused by different shooting positions and ensure the same motion trajectory, inspired by  [9, 25], we use a beam splitter to split the incident light into two beams with a ratio of 1:1 and enter the two camera lenses respectively, as shown in Fig. 2. In order to receive the natural incident light from the same viewpoint, we design and print an opaque 3D model box, which is also used to hold the beam splitter and connect the cameras. These components are concentrated on an optical breadboard and then fixed on the tripod to improve stability.
@@ 2.2 Data Collection
We use two Canon EOS R10 cameras to collect paired videos. Specifically,to minimize the loss of detail in shadows and highlights, we set one camera to automatic exposure mode and capture video in Canon Log [^1] https://cam.start.canon/en/C004/manual/html/UG-03_Shooting-2_0080.html format, which retains more visual information throughout the dynamic range. We then hand it over to professional photographers using DaVinci Resolve Studio [^2] https://www.blackmagicdesign.com/products/davinciresolve/studio software for manual rendering, ultimately producing high-quality and normal-exposure videos. The other camera captures low-quality abnormal exposure videos with the commonly used sRGB format. All the other settings are set to default values to simulate real capture scenarios. For each scene, we collect paired short 10-second videos at 50 FPS to capture both slow and fast motions.
@@ 2.3 Data Processing
Due to the slight errors in manually assembling the above hardware, there is still misalignment between the paired videos. Therefore, we utilize a two-stage frame alignment strategy to obtain aligned pairs. First, we estimate a homography matrix between the under-/overexposed frame and the corresponding normal exposed frame using the matched SIFT [18] key points. In this way, we can roughly crop matching regions in corresponding frames. Then we utilize a traditional flow estimation algorithm called DeepFlow [32] to perform pixel-wise alignment. Finally, we use center cropping to remove alignment artifacts around boundaries, producing precise spatial aligned frames. It should be noted that we only perform alignment correction on normal exposed frames and make the low-quality input of our network consistent with real captured abnormal exposed frames. In addition, we introduce lightness-order-error (LOE) [26] and optical flow [6] to quantitatively evaluate alignment performance and motion activity. As shown in Fig. 4, the paired frames of our dataset are better aligned and have faster movement than that of the low-light enhancement datasets including DRV [5] and SDSD [25], indicating stronger performance in real dynamic scenes.
We totally captured 119 groups of videos with a total of 21,951 frames to form our DIME dataset. Each video consists of 100-200 frames, and the resolution is 1,920 $\times$ 1,080. Fig. 3 presents some statistical analysis results. Fig. 5 gives three examples of our aligned video pairs under under-/over-exposure and normal conditions. Our captured videos vary from indoor to outdoor in real scenes and include camera and object motion types. See more detailed information in the supplementary file.
Figure 6: Overview of our framework. It contains three modules including Multi-frame Phase Alignment (MPA), Dual-stream Illumination Construction (DIC), and Two-stage Synthesis Restoration (TSR).
Figure 7: The detail of the Multi-frame Phase Alignment (MPA) module. We use discrete Fourier transformation to map the frames from pixel space to Fourier space, then adopt the phase branch for alignment.
@@ 3.1 Overview
The Video Exposure Correction (VEC) task can be formulated as seeking a mapping function $\mathcal{F}$, which maps consecutive 8-bit per channel sRGB frames $\mathcal{I}$ to enhanced frames $\mathcal{O}$ such that $\mathcal{O}=\mathcal{F}(\mathcal{I})$. During the training step, given 2N+1 consecutive frames $\mathcal{I}_{[t-N:t+N]}$ captured under different exposure conditions, the center frame is denoted as the reference frame while the others are supporting frames. The proposed end-to-end VECNet aims to restore the exposure of the reference frame $\mathcal{I}_{t}$ with the help of supporting frames, thereby generating the output frame $\mathcal{O}_{t}$. Based on Retinex theory [14], the overall architecture of VECNet is shown in Fig. 6, which is composed of three sub-networks: Multi-frame Phase Alignment Module (MPA), Dual-stream Illumination Construction Unit (DIC) and Two-stage Synthesis Restoration Unit (TSR). MPA module first aligns supporting frames to the reference frame to maintain temporal consistency and then learns the reflectance map together. Next, DIC unit estimates dual-path illumination maps for the reference frame to adaptively adjust its underexposure and overexposure factors. Finally, TSR unit fuses the above illumination and reflectance maps at feature and image levels, obtaining the high-quality output frame $\mathcal{O}_{t}$ with proper exposure.
3.2 Multi-frame Phase Alignment
Directly applying image exposure correction [1, 11] to each frame can cause flickering [25] due to ignoring temporal information. To take advantage of it, we take consecutive frames as the model’s input. However, since the texture information of different frames is misaligned, a direct fusion of multiple frames could cause blur and artifacts [22, 25]. Existing works mainly warp supporting frames to the reference frame by pyramid cascading and deformable convolution [28, 35] or group shift operations [16] directly at the feature level. However, in the video exposure correction task, There are differences in exposure between frames as camera motion produces different light positions. Such exposure mutation interferes with alignment. Recently,  [11] conducts exposure correction from a Fourier-based perspective, i.e., the amplitude component of an image reflects the lightness representation, while the phase component corresponds to structures and is less related to lightness. To this end, we take multi-frame alignment at the phase level in this section, as shown in Fig. 7.
Specifically, given a channel in single frame $\mathit{x}\in\mathbb{R}^{H\times W}$, the Fourier transform $\mathcal{F}$ converts $\mathit{x}$ to a complex component $\mathit{X}$ in Fourier space, which can be written as:
@@ 3.2 Multi-frame Phase Alignment
Then, amplitude $\mathcal{A}(\mathit{X})$ and phase $\mathcal{P}(\mathit{X})$ components can be divided from the complex component $\mathit{X}$:
@@
where $\mathit{R}$ and $\mathit{I}$ denote the real and imaginary parts of the complex $\mathit{X}$. We also apply the same strategy to each channel of input frames and then get consecutive amplitude components $\mathcal{A}(\mathit{X}_{[t-N:t+N]})$ and phase components $\mathcal{P}(\mathit{X}_{[t-N:t+N]})$. Since phase components respond more to the structure information and are less affected by the lightness while amplitude components are just the opposite, the phase components from supporting frames are warped to that of the reference frame by deformable convolution network (DCN). We compute the offset $\delta p^{k}_{t+i},i\in[-N,N],i\neq 0$ learned from corresponding $\mathcal{P}(\mathit{X}_{t+i})$ and $\mathcal{P}(\mathit{X}_{t})$ for the DCN. The aligned phase component $\tilde{\mathcal{P}}(\mathit{X}_{t})$ is obtained with the learned offset. The process can be formulated as:
$\begin{split}\delta p^{k}_{t+i}&=\mathcal{M}_{i}(\mathcal{P}(\mathit{X}_{t+i})\odot\mathcal{P}(\mathit{X}_{t})),\\ \tilde{\mathcal{P}}(\mathit{X}_{t+i})&=\mathcal{G}_{i}(DCN(\mathcal{P}(\mathit{X}_{t+i}),\delta p^{k}_{t+i})),\end{split}$ (3)
where $\odot$ denotes channel concatenation, $\mathcal{M}_{i}$ and $\mathcal{G}_{i}$ are the mapping function composed of several convolution layers. In addition, we obtain $\bar{\mathcal{A}}(X)$ aggregate multi-frame amplitude component to achieve unified representation. Finally, we recombine the aligned phase component and aggregated amplitude component by inverse Fourier transform $\mathcal{F}^{-1}$. The process is formulated as:
$\begin{split}\bar{\mathcal{A}}(X)&=\mathcal{H}[\mathcal{A}(\mathit{X}_{t-N})\odot...\mathcal{A}(\mathit{X}_{t+N})],\\ \tilde{x}_{t+i}&=\mathcal{F}^{-1}(\bar{\mathcal{A}}(X),\tilde{\mathcal{P}}(\mathit{X}_{t+i})),\end{split}$ (4)
where $\mathcal{H}$ is the aggregating function composed of several convolution and relu function layers. Then the aligned features consisting of multiple single-channel $\tilde{x}_{t+i}$ are sent to learn the reflectance map.

### v3

@@ Abstract
Exposure correction aims to enhance visual data suffering from improper exposures, which can greatly improve satisfactory visual effects. However, previous methods mainly focus on the image modality, and the video counterpart is less explored in the literature. Directly applying prior image-based methods to videos results in temporal incoherence with low visual quality. Through thorough investigation, we find that the development of relevant communities is limited by the absence of a benchmark dataset. Therefore, in this paper, we construct the first real-world paired video dataset, including both underexposure and overexposure dynamic scenes. To achieve spatial alignment, we utilize two DSLR cameras and a beam splitter to simultaneously capture improper and normal exposure videos. Additionally, we propose an end-to-end video exposure correction network, in which a dual-stream module is designed to deal with both underexposure and overexposure factors, enhancing the illumination based on Retinex theory. The extensive experiments based on various metrics and user studies demonstrate the significance of our dataset and the effectiveness of our method. The code and dataset are available at https://github.com/kravrolens/VECNet.
@@ 1 Introduction
Figure 1: The above sub-figure is the benchmark for our proposed video exposure correction, while the below sub-figures are the visual comparison between an image-based method LACT [2] and our proposed method. LACT takes single frames as input and results in temporal exposure incoherence with low visual quality. Our method utilizes temporal information to achieve consecutive exposures.
The images or videos taken across various scenarios may not always be ideal due to the changeable lighting conditions, which face underexposure and overexposure problems and yield unsatisfactory visual effects. How to improve inappropriately exposed visual data is gradually attracting the attention of researchers in the multimedia community. Due to the diversity of scenes and lighting conditions, the exposure correction process of visual data is particularly complicated, and overexposure and underexposure may exist at the same time (mixed exposure), which further enlarges the discrepancies of operations. Thus, many exposure correction methods have been proposed to handle this challenge problem automatically by utilizing deep learning technologies. Correcting underexposure and overexposure to normal exposures are much different from each other, which leads to large discrepancies in training an image-based multi-exposure correction network. This challenge is further compounded when dealing with videos captured in dynamic scenes.
In recent years, fueled by various paired image exposure dataset [3, 1], several methods [13, 12, 32, 26, 36] that focus on learning image exposure correction have been proposed. However, these methods generally do not handle videos of dynamic scenes, as directly applying them in a frame-by-frame manner often results in inter-frame exposure inconsistency with low visual quality [20, 16], as shown in the below sub-figures of Fig. 1. Some methods [34, 28, 9] focus on enhancing underexposed videos taken from low-light dynamic scenes. However, these methods can only achieve a single exposure correction. Due to the discrepant representations between underexposure and overexposure, the correction procedures differ greatly from each other, as shown in Fig. 3(a). Training a mixture of multi-exposure data in their models often leads to poor performance across different exposure levels, rendering these methods unsuitable for practical applications.
@@
To overcome the above challenges, in this work, we focus on the multi-exposure correction presented in the real-world videos for the first time. As of our current understanding, there is a noticeable absence of well-studied works or benchmark datasets in this domain. On the one hand, we aim to collect a high-quality video dataset with both underexposure and overexposure, and the main difficulties are as follows. First, collecting paired video datasets requires capturing two videos, one under abnormal exposure conditions and another under normal light of the same dynamic scene and camera motion. Second, it requires accurate alignment of each pair of corresponding frames in the two videos in both spatial and temporal dimensions. To address the above issues, we first design a two-camera system with a beam splitter to ensure no parallax between the two cameras with different aperture and ISO parameter settings. Two DSLR cameras simultaneously capture improper and normal exposure videos. Then we perform precise alignment on the captured pairs to make them aligned with each other. Furthermore, the well-exposed videos are manually rendered by professional photographers and then serve as the ground truth of the dataset. Finally, we construct a new dataset of 119 high-quality videos named DIME, which stands for ”Dynamic scenes In Multiple Exposure”. It contains diverse real-world scenes, camera and object motions, and each paired video is accurately aligned along the spatial and temporal dimensions.
On the other hand, since there are varying degrees of improper exposure in the constructed dataset, existing methods prove inadequate for deployment. Thus, on the other hand, we propose an end-to-end Video Exposure Correction Network (VECNet) to enhance the videos with multiple and improper exposures. Specifically, to adaptively learn the overexposed and underexposed representations, we formulate a dual-stream strategy to enhance various illumination components based on Retinex theory. For complementary reflectance component learning, we design a novel multi-frame alignment module that aligns neighboring frames into the middle one to effectively obtain comprehensive feature representations. In brief, our contributions can be summarized as follows:
@@
(1) We construct the first high-quality paired video exposure correction dataset for dynamic real-world scenes with multiple exposures, camera and object motions, and precise spatial alignment.
@@
(2) We propose an effective exposure correction network based on Retinex theory to enhance overexposed and underexposed videos.
(3) We conduct extensive experiments to demonstrate the superiority of our dataset and method.
@@ 2 Our DIME Dataset
Figure 2: We built the optical system to capture the normal and over-/under- exposure video pairs.
Figure 3: Statistics on our DIME dataset. (a) The input-ground truth brightness mapping curve statistics. (b) Luminance distribution for overexposure videos. (c) Luminance distribution for underexposure videos. (d) Luminance distribution for ground truth videos.
To learn video exposure correction tasks by training our model, we need a large number of videos, including realistic overexposure and underexposure errors and corresponding paired and properly exposed ground truth videos. As discussed in Sec. 2, such datasets are currently not publicly available to support video exposure correction research. Therefore, our first task is to create a novel dataset that captures various videos in dynamic and real-world scenes.
@@ 2.1 Hardware Design
Capturing multi-exposure image and video pairs for static scenes can be easily realized with a single-exposure adjustable DSLR camera [9] or an electric slide rail [28], which make it impractical to capture paired videos in wider dynamic scenes. We consider utilizing two cameras with different apertures and ISO parameter settings. To keep the cameras in sync, we use an infrared remote control to signal two camera-mounted receivers for simultaneous capture. To prevent parallax problems caused by different shooting positions and ensure the same motion trajectory, inspired by  [10, 28], we use a beam splitter to split the incident light into two beams with a ratio of 1:1 and enter the two camera lenses, respectively, as shown in Fig. 2. In order to receive the natural incident light from the same viewpoint, we design and print an opaque 3D model box, which is also used to hold the beam splitter and connect the cameras. These components are concentrated on an optical breadboard and then fixed to the tripod to improve stability.
@@ 2.2 Data Collection
We use two Canon EOS R10 cameras to collect paired videos. Specifically, to minimize the loss of detail in shadows and highlights, we set one camera to automatic exposure mode and capture video in Canon Log [^1] https://cam.start.canon/en/C004/manual/html/UG-03_Shooting-2_0080.html format, which retains more visual information throughout the dynamic range. Then we hand it over to professional photographers using DaVinci Resolve Studio [^2] https://www.blackmagicdesign.com/products/davinciresolve/studio software for manual rendering, ultimately producing high-quality and normal-exposure videos. The other camera captures low-quality abnormal exposure videos with the commonly used sRGB format. All the other settings are set to default values to simulate real capture scenarios. We collect paired 10-second videos for each dynamic scene.
@@ 2.3 Data Processing
Due to the slight errors in manually assembling the above hardware, there is still a misalignment between the paired videos. Therefore, we utilize a two-stage frame alignment strategy to obtain aligned pairs. First, we estimate a homography matrix between the under-/overexposed frame and the corresponding normal exposed frame using the matched SIFT [19] key points. In this way, we can roughly crop matching regions into corresponding frames. Then we utilize a traditional flow estimation algorithm called DeepFlow [35] to perform pixel-wise alignment. Finally, we use center cropping to remove alignment artifacts around boundaries, producing precise spatially aligned frames. Note that we only perform alignment correction on normal exposed frames and make the low-quality input of our network consistent with real captured abnormal exposed frames. Additionally, we introduce lightness-order-error (LOE) [29] and optical flow [7] to quantitatively evaluate alignment performance and motion activity. As shown in Fig. 4, the paired frames of our dataset are better aligned and have larger movement than that of the low-light enhancement datasets, including DRV [5] and SDSD [28], indicating stronger performance in real dynamic scenes.
We divide each video into multiple continuous, non-overlapping 200-frame clips. We totally captured 119 groups of clips with a total of 21,951 frames to form our DIME dataset. And the resolution is 1,920$\times$1,080. Fig. 3 presents some statistical analysis results. Fig. 5 gives three examples of our aligned video pairs under under-/over-exposure and normal conditions. Our captured videos vary from indoor to outdoor in real scenes and include camera and object motion types.
Figure 6: Overview of our framework. It contains three modules, including Multi-frame Fourier Alignment (MFA), Dual-stream Illumination Construction (DIC), and Two-stage Synthesis Restoration (TSR).
@@ 3.1 Overview
The Video Exposure Correction (VEC) task can be formulated as seeking a mapping function $\mathcal{F}$, which maps consecutive 8-bit per channel sRGB frames $\mathcal{I}$ to enhanced frames $\mathcal{O}$ such that $\mathcal{O}=\mathcal{F}(\mathcal{I})$. During the training step, given $12N+$ 1 consecutive frames $\mathcal{I}_{[t-N:t+N]}$ captured under different exposure conditions, the center frame is denoted as the reference frame while the others are supporting frames. The proposed end-to-end VECNet aims to restore the exposure of the reference frame $\mathcal{I}_{t}$ with the help of supporting frames, thereby generating the output frame $\mathcal{O}_{t}$. Based on Retinex theory [15], the overall architecture of VECNet is shown in Fig. 6, which is composed of three sub-networks: Multi-frame Fourier Alignment Module (MFA), Dual-stream Illumination Construction Unit (DIC) and Two-stage Synthesis Restoration Unit (TSR). The MFA module first aligns supporting frames with the reference frame to maintain temporal consistency and then learns the reflectance map together. Next, the DIC unit estimates dual-path illumination maps for the reference frame to adaptively adjust its underexposure and overexposure factors. Finally, the TSR unit fuses the above illumination and reflectance maps at feature and image levels, obtaining the high-quality output frame $\mathcal{O}_{t}$ with proper exposure.
3.2 Multi-frame Fourier Alignment
Directly applying image exposure correction [1, 12] to each frame can cause flickering [28] due to ignoring temporal information. To take advantage of the multi-frame information, we take consecutive frames as the model’s input. However, since the texture information of different frames is misaligned, a direct fusion of multiple frames could cause blur and artifacts [23, 28]. Existing works mainly warp supporting frames to the reference frame by pyramid cascading and deformable convolution [31, 37] or group shift operations [17] directly at the feature level. However, in the video exposure correction task, there are differences in degraded exposure between frames as camera motion produces different light positions. Such exposure mutations interfere with alignment. Recently,  [12] conducts exposure correction from a Fourier-based perspective, i.e., the amplitude component of an image reflects the lightness representation, while the phase component corresponds to structures and is less related to lightness. Therefore, we reduce the influence of lightness factors on alignment by adapting the amplitude component. To this end, we take multi-frame alignment using the Fourier transform in this section, as shown in Fig. 7.
Specifically, given a channel in a single frame $\mathit{x}\in\mathbb{R}^{H\times W}$, the Fourier transform $\mathcal{F}$ converts $\mathit{x}$ to a complex component $\mathit{X}$ in Fourier space, which can be written as:
@@ 3.2 Multi-frame Fourier Alignment
Then, the amplitude $\mathcal{A}(\mathit{X})$ and phase $\mathcal{P}(\mathit{X})$ components can be divided from the complex component $\mathit{X}$:
@@
where $\mathit{R}$ and $\mathit{I}$ denote the real and imaginary parts of the complex $\mathit{X}$. We also apply the same strategy to each channel of input frames and then get consecutive amplitude components $\mathcal{A}(\mathit{X}_{[t-N:t+N]})$ and phase components $\mathcal{P}(\mathit{X}_{[t-N:t+N]})$. Since the amplitude component responds more to the lightness, we obtain $\bar{\mathcal{A}}(X)$ by aggregating multi-frame amplitude to achieve a unified exposure representation:
$\begin{split}\bar{\mathcal{A}}(X)&=\mathcal{H}[\mathcal{A}(\mathit{X}_{t-N})\odot...\mathcal{A}(\mathit{X}_{t+N})],\\ \end{split}$ (3)
where $\mathcal{H}$ is the aggregating function composed of several convolution and relu function layers. Then the phase component $\tilde{\mathcal{P}}(X_{t+i})$ of the supporting frames and aggregated amplitude component $\bar{\mathcal{A}}(X)$ are recombined by the inverse Fourier transform $\mathcal{F}^{-1}$. The process is formulated as:
$\begin{split}\tilde{x}_{t+i}&=\mathcal{F}^{-1}(\bar{\mathcal{A}}(X),\tilde{\mathcal{P}}(\mathit{X}_{t+i})),\end{split}$ (4)
Then the supporting frames are warped to those of the reference frame by the deformable convolution network (DCN) [6]. We compute the offset $\delta x_{t+i},i\in[-N,N],i\neq 0$ learned from corresponding $\tilde{x}_{t+i})$ and $\tilde{x}_{t})$ for the DCN. The aligned phase component $\tilde{\mathcal{P}}(Y_{t})$ is obtained with the learned offset. The process can be formulated as:
$\begin{split}\delta x_{t+i}=\mathcal{M}_{i}(\tilde{x}_{t+i}\odot\tilde{x}_{t}),\quad Y_{t+i}=\mathcal{G}_{i}(DCN(\tilde{x}_{t+i},\delta x_{t+i})),\end{split}$ (5)
where $\odot$ denotes channel concatenation, and $\mathcal{M}_{i}$ and $\mathcal{G}_{i}$ are the mapping functions composed of several convolution layers. Then the aligned features consisting of multiple single-channel $Y_{t+i}$ are sent to learn the reflectance map.
Figure 7: The details of the Multi-frame Fourier Alignment module. We apply the discrete Fourier transformation to map the frames from pixel space to Fourier space, then adopt the deformable convolution for alignment.